OpenStudy (anonymous):

solve by elimination. 5x+6y=1 10x+12y=2 need help to make sure i did this right. thanks

OpenStudy (anonymous):

infinite solutions

OpenStudy (anonymous):

so how is the graph supped to look?

OpenStudy (anonymous):

same line

OpenStudy (anonymous):

minusdouble the 1st or minushalve the 2nd. 0=0

OpenStudy (anonymous):

one line

OpenStudy (anonymous):

so how about this one. 8x-9y=-63 9y-8x=63

OpenStudy (anonymous):

what's the ordered pair and the graph? i'm lost here

OpenStudy (anonymous):

8x-9y=-63 -8x+9y=63 infinite solutions one line

OpenStudy (anonymous):

thanks